40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Jun 17 11:05:42 2010, Frank, Lab Infrastructure, HVAC, HVAC now working again 
    Reply  Fri Jun 18 00:48:11 2010, rana, Lab Infrastructure, HVAC, HVAC now working again 
Message ID: 817     Entry time: Thu Jun 17 11:05:42 2010     Reply to this: 819
Author: Frank 
Type: Lab Infrastructure 
Category: HVAC 
Subject: HVAC now working again 

a minute ago a guy showed up in the lab and wanted to talk to Aidan but couldn't find him. He told me that the air conditioning is now working again and i should let him know.

Particle count in the ATF was about 1.5million before, now decreasing and already below 100k. Can't post data as the usb-to-serial converter is still used at 40m...

ELOG V3.1.3-